# Measuring Human and AI Values based on Generative Psychometrics with Large Language Models

### v1

@@ Abstract
Human values and their measurement are long-standing interdisciplinary inquiry. Recent advances in AI have sparked renewed interest in this area, with large language models (LLMs) emerging as both tools and subjects of value measurement. This work introduces Generative Psychometrics for Values (GPV), an LLM-based, data-driven value measurement paradigm, theoretically grounded in text-revealed selective perceptions. We begin by fine-tuning an LLM for accurate perception-level value measurement and verifying the capability of LLMs to parse texts into perceptions, forming the core of the GPV pipeline. Applying GPV to human-authored blogs, we demonstrate its stability, validity, and superiority over prior psychological tools. Then, extending GPV to LLM value measurement, we advance the current art with 1) a psychometric methodology that measures LLM values based on their scalable and free-form outputs, enabling context-specific measurement; 2) a comparative analysis of measurement paradigms, indicating response biases of prior methods; and 3) an attempt to bridge LLM values and their safety, revealing the predictive power of different value systems and the impacts of various values on LLM safety. Through interdisciplinary efforts, we aim to leverage AI for next-generation psychometrics and psychometrics for value-aligned AI.  [^1] Our code is available at https://github.com/Value4AI/gpv.
@@ 1 Introduction
Value theory, a cornerstone of philosophical inquiry, is paramount in guiding ethical decision-making and shaping societal norms [66]. Value measurement is a long-standing interdisciplinary endeavor for elucidating how specific values underpin and justify the worth of actions, objects, and concepts [76].
Traditional psychometrics often measure human values through self-report questionnaires, where participants rate the importance of various values in their lives. However, these tools are limited by response biases, resource demands, inaccuracies in capturing authentic behaviors, and inability to handle historical or subjective data [59]. Therefore, data-driven tools have been developed to infer values from textual data, such as social media posts [59, 26]. These tools can reveal personal values without relying on explicit self-reporting, but they are mostly dictionary-based, matching text to predefined value lexicons. Consequently, they often fail to grasp the nuances of semantic meaning and context-dependent value expressions. Additionally, these tools tend to be static and inflexible, relying on expert-defined lexicons that are not easily adaptable to new or evolving values.
The rise of large language models (LLMs) opens up new possibilities for data-driven value measurement. Models like GPT-4 have shown remarkable capabilities in understanding nuances in language and context. Recent studies have demonstrated that LLMs can effectively approximate annotators’ and even psychologists’ judgments on value-related tasks [86, 63]. This work introduces Generative Psychometrics for Values (GPV), an LLM-based, data-driven value measurement paradigm grounded in the theory of text-revealed selective perceptions. GPV overcomes the limitations of self-reports and dictionary-based tools by leveraging LLMs’ advanced semantic understanding. It extracts contextualized and value-laden perceptions (e.g., I believe that everyone deserves equal rights and opportunities.) from texts and decodes underlying values (e.g., Universalism) for arbitrary value systems. These perceptions function similarly to psychometric items in self-report questionnaires, often supporting or opposing specific values [76]. Notably, GPV enables the automatic generation of such items and their adaptation to any given data.
We begin by fine-tuning Llama 3 [24] for accurate perception-level value measurement. Benchmarked with established psychometric items, we demonstrate that the fine-tuned Llama 3 [24], referred to as ValueLlama, outperforms both state-of-the-art general and task-specific LLMs. Subsequently, we verify the capability of LLMs to parse texts into perceptions. The integration of perception-level value measurement and perception parsing forms the core pipeline of the GPV. Applying GPV to a large collection of human-authored blogs, we validate its stability and validity in measuring individual values, and its superiority over prior psychological tools.
@@
  • We introduce Generative Psychometrics for Values (GPV), a novel LLM-based value measurement paradigm grounded in text-revealed selective perceptions. GPV is accompanied by a fine-tuned LLM, ValueLlama, with state-of-the-art performance in perception-level value measurement (§ 3).
@@
  • Applying GPV to LLMs, we enable LLM value measurements based on their scalable, free-form, and context-specific outputs. With extensive evaluations across 17 LLMs and 4 value theories, we demonstrate the superiority of GPV and novel findings regarding value systems and LLM values (§ 5).
Figure 1: Illustrations of three measurement paradigms. (a) Self-reports require individuals to rate their agreement with expert-defined perceptions within a limited value space. (b) Dictionary-based methods count value-related lexicons given text data. The lexicons are expert-defined in a limited value space. (c) GPV automatically and dynamically generates perceptions from text data and learns to measure values in an open-ended value space.
@@ 2.1 Value Measurements for Human
The measurement of individual values is pivotal in elucidating the driving forces and mechanisms underlying human behavior [76, 66, 66]. Due to the intricate relationship between behavior and values, researchers have developed different measurement methods, including self-report questionnaires [81, 51, 87], behavioral observation [49, 52, 25, 64], and experimental techniques [70, 95, 56, 8]. Self-report methods involve participants themselves assessing their agreement with descriptions [48, 70] or ranking the importance of items [66]. Behavioral observation methods require experts to analyze how personal values manifest in real-life actions [9, 79]. Furthermore, experimental methods employ structured scenarios to isolate and analyze variables affecting human behavior [11, 92, 65]. While demonstrated effective in revealing how values guide decisions and interactions, these methods are hindered by response biases, resource demands, inaccuracies in capturing authentic behaviors, and inability to handle historical or subjective data [59, 14, 10].
Meanwhile, data-driven measurement tools have been developed to address the adverse effects of resource costs, external interference, and response biases. Dictionary-based tools such as LIWC dictionary [29] and personal values dictionary (PVD) [59] determine values by analyzing the frequency of value-related lexicons, flawed for overlooking nuanced semantics and contexts. Recent efforts to train deep learning models for value identification have largely focused on Schwartz’s values and are not validated for individual-level measurements [61, 55, 97, 86, 98]. [96, 42, 45] transform self-report inventories into interactive assessments based on LLMs, yet inherit many of the limitations of self-reports, such as the inability to handle historical or subjective data.
@@ 2.2 Value Measurements for LLMs
However, researchers have observed discrepancies between constrained and free-form LLM responses, and the latter is considered more practically relevant [93, 67, 63, 91]. The variability in LLM responses to subtle contextual changes also necessitates scalable and context-specific evaluation methods [40, 67, 98].
3 Generative Psychometrics for Values
3.1 Value Measurement with Selective Perceptions
Values are an individual’s concepts of transituational goals, reflecting interests within motivational domains and guiding principles in life [76]. Value measurement quantitatively evaluates the significance attributed to various values through individuals’ behavioral and linguistic data [3, 53, 66]. Given any pluralistic value system as a reference frame, we formalize the value measurement task as follows.
@@ 3.1 Value Measurement with Selective Perceptions
Here, $V=\{v_{1},v_{2},\ldots,v_{n}\}$ denotes a value system, where each $v_{i}$ represents a particular value dimension that motivates specific behaviors; $D$ denotes the individuals’ behavioral and linguistic data; and $\mathbf{w}=(w_{1},w_{2},\ldots,w_{n})$ is a value vector with $w_{i}$ indicating the relative importance of $v_{i}$.
Extensive research in psychology explores the underlying mechanisms of $f$, by which human values influence behaviors and behaviors reflect underlying values [3, 53, 76, 66]. GPV instantiates this process through value-driven selective perceptions [60, 4].
Personal values are demonstrable determinants of what individuals select to perceive, observe, and prioritize in their environment [60, 4]. For example, when considering a construction project of a new park, individuals who value Hedonism will emphasize the recreational benefits, while those who prioritize Economic Efficiency will focus on the project’s cost. These differing perceptions encode value-laden information and value orientations of individuals. Traditional psychometric inventories [77, 41, 6] adopt such contextual perceptions, commonly referred to as items, as organized stimulus. For example, VSM [31] include perceptions like "It is important to get recognition for good performance" and "It is important to have security of employment". They are designed to measure the value dimensions of Masculinity and Individualism, respectively. Each perception either supports or opposes a value. In the above examples, the former supports Masculinity, while the latter opposes Individualism. These traditional psychometric inventories mostly compile static and unscalable perceptions, covering a limited measurement range. They also necessitate an additional self-report process to assess the individual’s agreement with the items.
GPV, illustrated in Fig. 1, advances psychometrics by leveraging LLMs to dynamically generate perceptions according to given behavioral and linguistic data. Compared with traditional tools, GPV (1) effectively mitigates response bias and resource demands by dispensing with self-reports; (2) captures authentic behaviors instead of relying on forced ratings; (3) can handle historical or subjective data; (4) measures values in open-ended value spaces and easily adapts to new or evolving values without expert effort; and (5) enables more scalable and flexible value measurement.
@@ 3.2 Perception-level Value Measurement
We define perceptions, the atomic measurement units, by the following properties: (1) A perception should be value-laden and accurately describe the measurement subject, ensuring meaningful measurement. (2) A perception is an atomic measurement unit, ensuring unambiguous measurement. (3) A perception is well-contextualized and self-contained, ensuring that it alone is sufficient for value measurement. (4) All perceptions comprehensively cover all value-laden aspects of the measured subject, ensuring that no related content in the data is left unmeasured.
@@
We fine-tune Llama-3-8B [24] to perform perception-level value measurement in an open-ended value space, i.e., not restricted to a predefined value system. The fine-tuning involves the following two tasks [86] using datasets of ValueBench [63] and ValuePrism [86]: (1) Relevance classification determines whether a perception is relevant to a value. (2) Valence classification determines whether a perception supports, opposes, or remains neutral (context-dependent) towards a value. Both tasks are formulated as generating a label given a value and a perception. We present further training details in Appendix A.
@@
We refer to the fine-tuned Llama-3-8B as ValueLlama. Given a value system $V=\{v_{1},v_{2},\ldots,v_{n}\}$ and a sentence of perception $s$, we employ ValueLlama to calculate the relevance and valence probability distribution of each value $v_{i}$ to $s$, respectively denoted as $p_{rel}(\cdot|v_{i},s)$ and $p_{val}(\cdot|v_{i},s)$. Then, we compute $w_{i}$ using $p_{val}(\text{support}|v_{i},s)-p_{val}(\text{oppose}|v_{i},s)$ if the value is relevant ($p_{rel}(\cdot|v_{i},s)>0.5$) and its valence is classified as "support" or "oppose". Otherwise, $w_{i}$ is considered unmeasured.
@@
To evaluate the accuracy of perception-level value measurements, we hold out 50 values and 200 associated items (146 with "Supports" valence and 54 with "Opposes" valence) from ValueBench as test datasets, also ensuring the values are not included in ValuePrism. We use Kaleido [86] and GPT-4 Turbo [2] as baselines and present the comparison results in Table 1, showing that ValueLlama outperforms state-of-the-art general and task-specific LLMs in perception-level value measurement.
@@ 3.3 Parsing and Aggregation
To measure values at the individual level, we first parse the individual’s textual data into perceptions and then aggregate these perception-level measurement results.
Parsing Text into Perceptions.
We begin by recursively chunking the text into segments of 250 tokens. Next, we prompt GPT-3.5 Turbo to enumerate perceptions for each segment, guided by a background on human values, definitions of perceptions, and few-shot examples. The process is detailed in § B.1.
Evaluating Parsing Results.
The parsing results are deemed high-quality by specifically trained human annotators. The evaluation is detailed in § B.2.
Aggregation.
Individual-level measurements are obtained by averaging the perception-level measurements for each value [83].
@@ 4 GPV for Humans
This section measures human values using 791 blogs from the Blog Authorship Corpus [75]. We evaluate GPV through standard Psychological metrics including stability, construct validity, concurrent validity, and predictive validity, and demonstrate its superiority over established psychological tools.
@@ 4.1 Validation
As values are relatively stable psychological constructs for humans [69, 71, 39], we expect that the same individual should exhibit consistent value tendencies across different scenarios. Across 48888 perception-value pairs, 86.6% of the perception-level measurement results are consistent with the individual-level aggregated results, indicating desirable stability. Detailed results and extended discussions are shown in § C.1.
@@
Construct validity is the extent to which a test measures what it claims to measure. In Schwartz’s value system, some values are theoretically positively correlated, such as Self-Direction and Stimulation, while others are negatively correlated, such as Power and Benevolence. Altogether, the 10 Schwartz values form a circumplex structure [78, 82], where values that are closer together are more compatible, while those that are farther apart are more conflicting (Fig. 2(a)). We employ multidimensional scaling (MDS) to evaluate the construct validity of the Schwartz’s value system [22, 12]. Based on the correlation between values derived from GPV measurement results, we project both the 10 basic values and the 4 higher-order values onto two-dimensional MDS plots. Then, we assess whether their relative positions align with the theoretical structure. As illustrated in Fig. 2, basic values belonging to the same category (represented by the same color) generally cluster together. Higher-order opposing values are positioned farther apart. The relative positions of a few values do not strictly follow the theoretical structure. For example, Conservation is relatively distant from the other three higher-order values. Such deviations may reflect a gap between the values manifested by self-report and objective data [59]. Overall, the relative positioning of most values resembles the theoretically expected pattern in Fig. 2(a), indicating desirable construct validity.
@@
Concurrent validity is the extent to which a test correlates with other measures of the same construct administered simultaneously. Theoretically expected correlations can validate newly developed instruments [90, 47]. We evaluate the concurrent validity of GPV by comparing it with the personal values dictionary (PVD) [59, 26], a well-established measurement tool with proven reliability and validity. We analyze the correlations between GPV and PVD measurements, with the results of low-level values presented in § C.3 and high-level aggregated values in Table 2. The results indicate that among the 10 basic values, both identical values (e.g., SE-SE) and most compatible values (e.g., CO-SE) show positive correlations. In contrast, most opposing values (e.g., BE-AC) exhibit negative correlations. Similarly, within the 4 higher-order values, positive correlations are observed among identical values, whereas most opposing values display negative correlations. These correlations, though not strong, are theoretically expected, which supports the concurrent validity of GPV. § 4.2 discusses the cases where GPV misaligns with PVD.
@@
Predictive validity is the extent to which a test predicts future behavior or outcomes. We assess predictive validity by examining if our measurement results align with the blog authors’ gender-related socio-demographic traits. Previous research indicates that men statistically prioritize power, stimulation, hedonism, achievement, and self-direction, while women emphasize benevolence and universalism [80]. Our measurement results, presented in Table 3, reveal that men and women score higher on the values they typically prioritize, confirming the consistency of our measurements with established statistical theories.

### v2

@@ Abstract
Human values and their measurement are long-standing interdisciplinary inquiry. Recent advances in AI have sparked renewed interest in this area, with large language models (LLMs) emerging as both tools and subjects of value measurement. This work introduces Generative Psychometrics for Values (GPV), an LLM-based, data-driven value measurement paradigm, theoretically grounded in text-revealed selective perceptions. The core idea is to dynamically parse unstructured texts into perceptions akin to static stimuli in traditional psychometrics, measure the value orientations they reveal, and aggregate the results. Applying GPV to human-authored blogs, we demonstrate its stability, validity, and superiority over prior psychological tools. Then, extending GPV to LLM value measurement, we advance the current art with 1) a psychometric methodology that measures LLM values based on their scalable and free-form outputs, enabling context-specific measurement; 2) a comparative analysis of measurement paradigms, indicating response biases of prior methods; and 3) an attempt to bridge LLM values and their safety, revealing the predictive power of different value systems and the impacts of various values on LLM safety. Through interdisciplinary efforts, we aim to leverage AI for next-generation psychometrics and psychometrics for value-aligned AI.  [^1] Our code is available at https://github.com/Value4AI/gpv.
@@ 1 Introduction
Human values, a cornerstone of philosophical inquiry, are the fundamental guiding principles behind individual and collective decision-making [68, 73]. Value measurement is a long-standing interdisciplinary endeavor for elucidating how specific values underpin and justify the worth of actions, objects, and concepts [78, 36, 95, 42].
Traditional psychometrics often measure human values through self-report questionnaires, where participants rate the importance of various values in their lives. However, these tools are limited by response biases, resource demands, inaccuracies in capturing authentic behaviors, and inability to handle historical or subjective data [61]. Therefore, data-driven tools have been developed to infer values from textual data, such as social media posts [84, 61, 26]. These tools can reveal personal values without relying on explicit self-reporting, but they are mostly dictionary-based, matching text to predefined value lexicons. Consequently, they often fail to grasp the nuanced semantics and context-dependent value expressions. Additionally, these tools are inherently static and inflexible, relying on expert-defined lexicons that are not easily adaptable to new or evolving values.
The rise of large language models (LLMs), with their remarkable ability to understand semantic nuances, presents new possibilities for data-driven value measurement. Recent studies have demonstrated that LLMs can effectively approximate annotators’ and even psychologists’ judgments on value-related tasks [87, 65]. Building on these advancements, this work introduces Generative Psychometrics for Values (GPV), an LLM-based, data-driven value measurement paradigm grounded in the theory of text-revealed selective perceptions [62, 4, 84]. Perceptions are the way individuals interpret and evaluate the world around them, and are servants of interests, needs, and, values [62]. Such perceptions are revealed in self-expressing texts, such as blog posts, and are utilized as atomic value measurement units in GPV. The core idea of GPV is to extract contextualized and value-laden perceptions (e.g., "I believe that everyone deserves equal rights and opportunities.") from unstructured texts, decode underlying values (e.g., Universalism) for arbitrary value systems, and aggregate the results to measure individual values.
The perceptions in GPV function similarly to the static psychometric items (stimuli) in self-report questionnaires, which support or oppose specific values [78]. Notably, GPV enables the automatic generation of such items and their adaptation to any given data, overcoming the limitations of traditional tools (Fig. 1). By applying GPV to a large collection of human-authored blogs, we evaluate GPV against psychometric standards. GPV demonstrates its stability and validity in measuring individual values, and its superiority over prior psychological tools.
@@
  • We introduce Generative Psychometrics for Values (GPV), a novel LLM-based value measurement paradigm grounded in text-revealed selective perceptions (§ 3).
@@
  • Applying GPV to LLMs, we enable LLM value measurements based on their scalable, free-form, and context-specific outputs. With extensive evaluations across 17 LLMs, 4 value theories, and 3 measurement tools, we illustrate the superiority of GPV and uncover novel insights regarding value systems and LLM values (§ 5).
Figure 1: Illustrations of the three measurement paradigms. (a) Self-reports require individuals to rate their agreement with expert-defined perceptions. (b) Dictionary-based methods count expert-defined and value-related lexicons given text data. (c) GPV automatically and dynamically extracts perceptions from text data and learns to measure open-vocabulary values.
@@ 2.1 Value Measurements for Human
The measurement of individual values is pivotal in elucidating the driving forces and mechanisms underlying human behavior [78, 68]. Due to the intricate relationship between behavior and values, researchers have developed different measurement methods, including self-report questionnaires [82, 53, 88], behavioral observation [51, 54, 25, 66], and experimental techniques [72, 96, 58, 7]. Self-report methods involve participants themselves assessing their agreement with descriptions [50, 72] or ranking the importance of items [68]. Behavioral observation methods require experts to analyze how personal values manifest in real-life actions [8, 80]. Furthermore, experimental methods employ structured scenarios to isolate and analyze variables affecting human behavior [10, 92, 67]. However, these methods are hindered by response biases, resource demands, inaccuracies in capturing authentic behaviors, and inability to handle historical or subjective data [61, 13, 9].
On the other hand, data-driven tools partially address the adverse effects of resource costs, external interference, and response biases. Among them, dictionary-based tools such as LIWC dictionary [30] and personal values dictionary (PVD) [61] analyze the frequency of value-related lexicons, flawed for overlooking nuanced semantics and contexts. Recent efforts to train deep learning models for value identification have largely focused on Schwartz’s values and are not validated for individual-level measurements [63, 57, 98, 87, 99]. Other works transform self-report inventories into interactive assessments based on LLMs [97, 44, 47], yet inherit many of the limitations of self-reports, such as the inability to handle historical or subjective data.
@@ 2.2 Value Measurements for LLMs
However, researchers have observed discrepancies between constrained and free-form LLM responses, and the latter is considered more practically relevant [93, 69, 65, 91]. The variability in LLM responses to subtle contextual changes also necessitates scalable and context-specific evaluation methods [43, 69, 99], which this work aims to address.
3 Generative Psychometrics for Values (GPV)
3.1 Value Measurement Based on Selective Perceptions
Values are broad motivational goals and guiding principles in life [78]. Value measurement quantitatively evaluates the significance attributed to various values through individuals’ behavioral and linguistic data [3, 55, 68]. Given any pluralistic value system as a reference frame, we formalize the value measurement task as follows.
@@ 3.1 Value Measurement Based on Selective Perceptions
Here, $V=\{v_{1},v_{2},\ldots,v_{n}\}$ denotes a value system, where each $v_{i}$ represents a particular value dimension; $D$ denotes the individuals’ behavioral and linguistic data; and $\mathbf{w}=(w_{1},w_{2},\ldots,w_{n})$ is a value vector with $w_{i}$ indicating the relative importance of $v_{i}$.
Extensive research explores the underlying mechanisms of $f$, by which human values drive behaviors and behaviors reflect values [3, 55, 78, 68]. Most related to this work, self-reports (Fig. 1(a)) instantiate $f$ by self-rating the agreement with expert-defined items; dictionary-based methods (Fig. 1(b)) instantiate $f$ by counting expert-defined and value-related lexicons. Both tools conduct value measurement in a limited value space (e.g. 10 Schwartz’s values define a limited 10-dimensional value space) and are inherently static and inflexible.
GPV Overview.
In contrast, GPV (Fig. 1(c)) instantiates $f$ through selective perceptions, a process of selecting stimuli from the environment based on an individual’s interests, needs, and values [62, 4]. For example, when considering a construction project of a new park, individuals who value Hedonism will emphasize the recreational benefits, while those who prioritize Economic Efficiency will focus on the project’s cost. These differing perceptions encode value orientations. GPV leverages LLMs to automatically parse self-expressing texts into such perceptions, trains an LLM for perception-level and open-vocabulary value measurement, and aggregates the results as individual values. We elaborate on the perception-level value measurement in § 3.2, then parsing and aggregation in § 3.3.
@@ 3.2 Perception-Level Value Measurement
Perceptions are selective stimuli from the environment that reflect an individual’s interests, needs, and values [62]. Here, perceptions are utilized as atomic measurement units, ideally capturing the following properties [29]: (1) A perception should be value-laden and accurately describe the measurement subject, ensuring meaningful measurement. (2) A perception is an atomic measurement unit, ensuring unambiguous measurement. (3) A perception is well-contextualized and self-contained, ensuring that it alone is sufficient for value measurement. (4) All perceptions comprehensively cover all value-laden aspects of the measured subject, ensuring that no related content in the data is left unmeasured.
@@
We fine-tune Llama-3-8B [24] for perception-level and open-vocabulary value measurement. Its fine-tuning involves the following two tasks [87] using datasets of ValueBench [65] and ValuePrism [87]: (1) Relevance classification determines whether a perception is relevant to a value. (2) Valence classification determines whether a perception supports, opposes, or remains neutral (context-dependent) towards a value. Both tasks are formulated as generating a label given a value and a perception. We present further training details in Appendix A.
@@
We refer to the fine-tuned Llama-3-8B as ValueLlama. Given a value system $V=\{v_{1},v_{2},\ldots,v_{n}\}$ and a sentence of perception $s$, we employ ValueLlama to calculate the relevance and valence probability distribution of each value $v_{i}$ to $s$, respectively denoted as $p_{rel}(\cdot|v_{i},s)$ and $p_{val}(\cdot|v_{i},s)$. Then, we define $w_{i}$ as $p_{val}(\text{support}|v_{i},s)-p_{val}(\text{oppose}|v_{i},s)$ if the value is relevant ($p_{rel}(\cdot|v_{i},s)>0.5$) and its valence is classified as "support" or "oppose". Otherwise, $w_{i}$ is considered unmeasured. The prompts for inference are listed in Appendix A.
@@
To evaluate the accuracy of perception-level value measurements, we hold out 50 values and 200 associated items (146 with "Supports" valence and 54 with "Opposes" valence) from ValueBench as a test dataset, also ensuring the test values are not included in ValuePrism. Using the same zero-shot prompt, we measure the relevance and valence of the test items with Kaleido [87], GPT-4 Turbo [2], and ValueLlama. Table 1 presents the comparison results, indicating that ValueLlama outperforms state-of-the-art general and task-specific LLMs in zero-shot perception-level value measurement.
@@ 3.3 Parsing and Aggregation
To measure values at the individual level, GPV chunks long texts (e.g., blog posts) into segments and prompts an LLM (this work used GPT-3.5 Turbo) to parse each segment into perceptions. Parsing is guided by the background on human values, definitions of perceptions, and few-shot examples (§ B.1.) Then, GPV performs perception-level value measurement (§ 3.2) for the parsing results. Individual-level measurements are calculated by averaging the perception-level measurements for each value [83].
Evaluating LLM Parsing.
The parsing results are considered high-quality by trained human annotators. On average, the annotators agree that the parsing results meet the defined four criteria in over 85% of cases, deeming them suitable for further value measurement. The evaluation is detailed in § B.2.
3.4 Discussion
Relation to Self-Reports.
The items organized in self-report inventories are essentially perceptions that support or oppose specific values [78]. Compared to GPV, these traditional psychometric inventories compile static and unscalable perceptions, covering a limited measurement range. They also necessitate an additional self-report process to assess the individual’s agreement with the items.
Relation to Dictionary-Based Methods.
Both GPV and dictionary-based methods share the fundamental principle that values are embedded in language [84], and they each measure values through text data. However, dictionary-based methods depend on predefined lexicons for closed-vocabulary values and are far less expressive than GPV in capturing semantic nuances. Further analysis is presented in § 4.2.
Advantages of GPV.
Compared with traditional tools, GPV 1) effectively mitigates response bias and resource demands by dispensing with self-reports; 2) captures authentic behaviors instead of relying on forced ratings; 3) can handle historical or subjective data; 4) measures values in open-ended value spaces and easily adapts to new or evolving values without expert effort; and 5) enables more scalable and flexible value measurement.
@@ 4 GPV for Humans
This section measures human values using 791 blogs from the Blog Authorship Corpus [77], selected after filtering out low-quality entries (§ C.1). We evaluate GPV using standard psychological metrics including stability, construct validity, concurrent validity, and predictive validity, and demonstrate its superiority over established psychological tools.
@@ 4.1 Validation
As values are relatively stable psychological constructs for humans [71, 73, 41], we expect that the same individual should exhibit consistent value tendencies across different scenarios. Across 48,888 perception-value pairs, 86.6% of the perception-level measurement results are consistent with the individual-level aggregated results, indicating desirable stability. Detailed results and extended discussions are shown in § C.2.
@@
Construct validity is the extent to which a test measures what it claims to measure. In Schwartz’s value system, some values are theoretically positively correlated, such as Self-Direction and Stimulation, while others are negatively correlated, such as Power and Benevolence. Altogether, the 10 Schwartz values form a circumplex structure [79], where values that are closer together are more compatible, while those that are farther apart are more conflicting (Fig. 2(a)). We employ multidimensional scaling (MDS) [21, 11] on the value correlations obtained by GPV, and project both the 10 basic values and the 4 higher-order values onto two-dimensional MDS plots. Then, we assess whether their relative positions align with the theoretical structure. As illustrated in Fig. 2, basic values of the same category (represented by the same color) generally cluster together. Higher-order opposing values are positioned farther apart. The relative positions of a few values do not strictly follow the theoretical structure. For example, Conservation is relatively distant from the other three higher-order values. Such deviations may reflect a gap between the values manifested by self-report and objective data [61]. Overall, the relative positioning of most values resembles the theoretically expected pattern in Fig. 2(a), indicating desirable construct validity. More experimental details are provided in § C.3.
@@
Concurrent validity is the extent to which a test correlates with other measures of the same construct administered simultaneously. Theoretically expected correlations can validate newly developed instruments [49]. We evaluate the concurrent validity of GPV by comparing it with the personal values dictionary (PVD) [61], a well-established measurement tool with proven reliability and validity. We analyze the correlations between GPV and PVD measurements, with the results of low-level values presented in § C.4 and high-level aggregated values in Table 2. The results indicate that among the 10 basic values, both identical values (e.g., SE-SE) and most compatible values (e.g., CO-SE) show positive correlations; most opposing values (e.g., BE-AC) exhibit negative correlations. Similarly, within the 4 higher-order values, positive correlations are observed when measuring identical values, whereas most opposing values display negative correlations. These correlations, though not strong, are theoretically expected, which supports the concurrent validity of GPV. § 4.2 exemplifies the cases where GPV misaligns with PVD.
@@
Predictive validity is the extent to which a test predicts future behavior or outcomes. We assess predictive validity by examining if our measurement results align with the blog authors’ gender-related socio-demographic traits. Previous research indicates that, in a statistical sense, men prioritize power, stimulation, hedonism, achievement, and self-direction, while women emphasize benevolence and universalism [81]. Our measurement results, presented in Table 3, reveal that men and women score higher on the values they typically prioritize, confirming the consistency of our measurements with established psychological findings.